News

From Tardis
Jump to: navigation, search

Infrequently updated - for a more realistic picture see Recent Changes.

24 August 2017

  • The new Tardis room is AT 7.03.

11 April 2017

  • The School of Informatics have identified a room that may be used as a dedicated server room for Tardis on level 7 of Appleton Tower.

10 Feb 2017

  • It has been agreed that the School of Informatics will find a (yet to be determined) location for the Tardis servers within Appleton Tower.

1 May 2015

  • Tardis is now living in the Informatics Forum Data Center. Access is supervised and only within business hours (Mon-Fri).

14 October 2013

  • We have a webserver again! The wiki is back, and is authenticated against LDAP. Yay!


10 July 2012

  • Mantis returns! After a 7-8 month break we have a webserver again. Unfortunately the old webserver data is still stuck on assorted (possibly corrupted) SCSI drives, if anyone wants to have a go at pulling the data off let us know!
  • NFS shares have moved to a new server Astrid which is a Xen DomU running on Ukairo a HP ProLiant with two RAID arrays inc. 400GB of RAID5 which will store all homedir/maildir and eventually webdirs (if we get them back).

29 September 2011

  • Tardis is now living in Room 5.06 in Appleton Tower. We have 24/7 (with student card) and a key to get into the actual room (blip2 currently holds it (and the rack keys)). We also share the room with of the Compsoc committee whom also have keys. See ATMove for more information. Also: all of the servers passed a PAT test (for what it's worth)

28 September 2011

  • Tardis is moving again to Room 5.06 in Appleton Tower. We should have more space, better access (24/7) and a key to get in! [tbc]. Moving will probably happen from about 1pm - any help welcome!

16 November 2010

  • IPv6 'AAAA' entries in DNS have been disabled due to 6to4 tunnel being broken (router address is now externally firewalled preventing the 6in4 packets arriving).

29 October 2010

  • Ace is no longer in service as the router - it has been replaced by Midnight. However, the router is still 'ace' in dns to make the transition easier.

14 September 2010

  • Ace and King are being recommissioned as routers by Amazoph routing will be switched to the new boxes at some point in the near future.
  • Hopefully these will configured in a HA setup at some point.

24 August 2010

  • Ace has been decommissioned - the router service is now running on WOTAN. We had some downtime as a result, however all is well now.
  • We still need a more permanent solution for the main router.

16 March 2010

  • All boxes are now working as intended after the move.

22 February 2010

  • Tardis was moved from Forest Hill to the Informatics Forum. There was significant apparent downtime due to the main xen host (WOTAN) freezing on boot (required a reinstall of the dom0).

29 September 2009

  • Tardis had a meeting to discuss the future i.e. moving out of Forest Hill and other infrastructure related things.

7 August 2009

  • Six Sun U1 boxes were installed in the rack (with correct mountings - thanks Pert!), two of them will run Tardis services (possibly NFS/backup) and the other four will replace the five beowulf cluster boxes which were decommissioned to make room in the rack. Currently none of the boxes are running (no OS etc) but they should be fairly shortly.
  • Mantis, the new webserver, is currently being commissioned and should be in service soon.

31 July 2009

  • Mantis, the new webserver was setup using some complicated adaptors to plug in into the serial console. It should start operations shortly.

5 June 2009

  • The Sun Fire 280r that is to serve as our webserver was installed in the rack. We didn't have time to set it up properly/network it however.
  • Downtime was scheduled so we could also re-arrange the cabling. All dual-PSU boxes (ex. Oracle) now have two power supplies, one from each of the distro units.

27 April 2009

  • Wiki has been given a makeover so it can be a front for Tardis. It's still a bit slow but that is the webserver's fault.

07 February 2009

  • We have completed the Move of Piper to the new box. It remains called Piper however, as that made the move simpler.

13:12, 11 July 2008 (BST)

  • The move to Forest Hill went well. One PSU died, one disk had some issues and there's a bit more work to do on the web server but, other than that, tardis is working once more.

19:32, 14 March 2007 (UTC)

  • We had some downtime today for for power testing. All is well.

14:10, 22 November 2006 (GMT)

21:29, 5 July 2006 (BST)

  • People are on holiday. New switch going in soon. Mostly, all is well.

23:58, 18 February 2006 (GMT)

  • Downtime didn't happen in the end, we were all somewhat relieved.
  • NFS Transition onto piper went well, aside from a mild duplication of mail. Next on the agenda is the decommissioning of Baker and Colin.

19:20, 7 February 2006 (GMT)

  • Downtime happening soon. Sadly in a non-derministic fashion. Some things are out of our control.
  • The wiki is now publicly readable, but not editable. If you would like an account, speak to seth. Maybe one day we'll get LDAP auth.

00:57, 18 January 2006 (GMT)